OpenStudy (anonymous):

Help please! :) What value of x solves the following proportion? 9/6=x/8 6(3/4) 11 10(1/2) 12

OpenStudy (owlcoffee):

\[\frac{ 9 }{ 6 }=\frac{ x }{ 8 }\] As the problem says, it's a proportion, so let's multiply in cross: \[9(8)=x(6)\] Just to convert in a simple ecuation, now let's multiply: \[72=6x\] And solving for x: \[x=\frac{ 72 }{ 6 }\rightarrow x=12\] And done :)

OpenStudy (anonymous):

Thank you so much!:)